OpenStudy (maddie82502):

combine like terms 9+18s-7s A-9+11s B-20 C-9-11s D-20 medal and fan if u help

OpenStudy (maddie82502):

@ganeshie8

OpenStudy (anonymous):

As 9+18s-7s =9+11s As 9+18s-7s=9+11s 9+18s-7 is Option "A".... Hope this helped u!!!!

OpenStudy (anonymous):

thanks for the medal :D

OpenStudy (maddie82502):

thank u @AskMeLotss and it helped a lot

OpenStudy (anonymous):

your welcome :D